TPC weekly
2021-06-09 15:00
America/New York
Wednesday, 9 June 2021
virtual , at 19:00 (GMT), duration : 00:00
virtual , at 19:00 (GMT), duration : 00:00
agenda:
- operation
- calibration status
- FF field
Yuri posted additional slide on FF DCA_y
https://drupal.star.bnl.gov/STAR/system/files/OO200GeV2021RFvsFF_0.pdf
Tpc meeting June 9, 2021
Present: Gene, Yuri, Dan, Babu, Alexei, HongWei, Flemming
Hardware:
Humidity has been creeping up a bit. Yesterday some high current in fieldcage due to humidity.
Quote from Gene
“I find that the blower A/C being off on June 5th did raise the inner field cage currents during two physics runs:
22156024 and 026
I propose simply to mark these two runs as bad. I would advise in situations like this to stop the fill and address the blower ASAP rather than waiting for RHIC operators to reach their own end-of-fill.
The number of auto-recovery has gone up during the 3.985 GeV FXT . Beam is likely dirtier. Date for 24:s4 looks fine
Operation:
The ops crew noted alarm for magnet current at ops meeting. It seems to be correlated to the lower mean current observed the last day. This was observed earlier in the run by Yuri see https://drupal.star.bnl.gov/STAR/blog/fisyak/averaged-star-magnet-current-table. The large oscillation is due to measurements (+-2A)– this is understood. The drop today as well as drops earlier in the run is not. The oscillation gives a K0S mass change of ~0.4 GeV/c**2.
Yuri was asked to look at mean of K0S mass for the periods (or a period) where the mean is down to -4112 for average to identify if it is due to measurement or to change in magnetic field.
Alexei has earlier noted that for next year we should have a NMR installed with high precision field measurement.
Calibration:
At this point the TPC calibration are done for run-19.
Babu plans to have an update next week for the run-20 iterations.
Yuri had an additional plot of the FF local DCA_y ivs. Sector number
https://drupal.star.bnl.gov/STAR/system/files/OO200GeV2021RFvsFF_0.pdf
The fits correspond to a rotation around the vertical axis of the field relative to TPC, and is of similar magnitude as the measured difference between field and TPC at zero-filed and incorporated in the ‘twist distortion correction. Yuri has discussed with Alexei and Bill C, to see if this can be understood i.e why changing the field to result in such change.
- operation
- calibration status
- FF field
Yuri posted additional slide on FF DCA_y
https://drupal.star.bnl.gov/STAR/system/files/OO200GeV2021RFvsFF_0.pdf
Tpc meeting June 9, 2021
Present: Gene, Yuri, Dan, Babu, Alexei, HongWei, Flemming
Hardware:
Humidity has been creeping up a bit. Yesterday some high current in fieldcage due to humidity.
Quote from Gene
“I find that the blower A/C being off on June 5th did raise the inner field cage currents during two physics runs:
22156024 and 026
I propose simply to mark these two runs as bad. I would advise in situations like this to stop the fill and address the blower ASAP rather than waiting for RHIC operators to reach their own end-of-fill.
The number of auto-recovery has gone up during the 3.985 GeV FXT . Beam is likely dirtier. Date for 24:s4 looks fine
Operation:
The ops crew noted alarm for magnet current at ops meeting. It seems to be correlated to the lower mean current observed the last day. This was observed earlier in the run by Yuri see https://drupal.star.bnl.gov/STAR/blog/fisyak/averaged-star-magnet-current-table. The large oscillation is due to measurements (+-2A)– this is understood. The drop today as well as drops earlier in the run is not. The oscillation gives a K0S mass change of ~0.4 GeV/c**2.
Yuri was asked to look at mean of K0S mass for the periods (or a period) where the mean is down to -4112 for average to identify if it is due to measurement or to change in magnetic field.
Alexei has earlier noted that for next year we should have a NMR installed with high precision field measurement.
Calibration:
At this point the TPC calibration are done for run-19.
Babu plans to have an update next week for the run-20 iterations.
Yuri had an additional plot of the FF local DCA_y ivs. Sector number
https://drupal.star.bnl.gov/STAR/system/files/OO200GeV2021RFvsFF_0.pdf
The fits correspond to a rotation around the vertical axis of the field relative to TPC, and is of similar magnitude as the measured difference between field and TPC at zero-filed and incorporated in the ‘twist distortion correction. Yuri has discussed with Alexei and Bill C, to see if this can be understood i.e why changing the field to result in such change.
OO Multiplicity
Updated on Wed, 2021-06-09 19:00. Originally created by slhuang on 2021-06-09 14:51.
OO 200 GeV 2021 RF vs FF
Updated on Wed, 2021-06-09 13:54. Originally created by fisyak on 2021-06-09 13:54.
HCal pi0 finding for data
Updated on Wed, 2021-06-09 11:01. Originally created by liangxl on 2021-06-09 11:01.Reference
Speaker : Xilin Liang
Talk time : 11:00, Duration : 00:20
HCal pi0 finding for data
Updated on Wed, 2021-06-09 11:00. Originally created by liangxl on 2021-06-09 11:00.Reference
Speaker : Xilin Liang
Talk time : 11:00, Duration : 00:20
Spin PWG Start-Less TOF And Jet Tree Code Update
Updated on Wed, 2021-06-09 07:41. Originally created by bassam on 2021-06-09 07:41.
Lambda global polarization in isobar( E/W EPD effect)
Updated on Wed, 2021-06-09 04:49. Originally created by xgou on 2021-06-09 04:49.
Study of the nuclear deformation in isobar collisions
Updated on Tue, 2021-06-08 20:49. Originally created by chunjian on 2021-06-08 19:38.for bulk-corr meeting
AOB
Updated on Tue, 2021-06-08 19:28. Originally created by dongx on 2021-06-08 19:28.Plan for 64-bit Test
Updated on Tue, 2021-06-08 19:28. Originally created by dongx on 2021-06-08 19:28.Reference
Speaker : Jason Webb, Rongrong Ma
Talk time : 12:30, Duration : 00:15
Input from the PWGC on the test samples.
CA tracker update
Updated on Wed, 2021-06-09 13:11 by kehw. Originally created by dongx on 2021-06-08 19:27.Run19 production readiness
Updated on Tue, 2021-06-08 19:24. Originally created by dongx on 2021-06-08 19:24.trigger upgrade status
Updated on Tue, 2021-06-08 09:41. Originally created by ruanlj on 2021-06-08 09:41.Speaker : Lijuan Ruan
Talk time : 09:30, Duration : 00:20
trigger upgrade status
Updated on Tue, 2021-06-08 09:41. Originally created by ruanlj on 2021-06-08 09:41.Speaker : Lijuan Ruan
Talk time : 09:30, Duration : 00:20
Embedding in 2012 pp vs data
Updated on Tue, 2021-06-08 09:21. Originally created by licenrob on 2021-06-08 09:21.Talk time : 09:00, Duration : 00:20
Added comparison to Petr's own production of data in SL13b_embed, this one should (and does) match Dmitry's own production in the same library.